### v1

# Hedgehog topological defects in 3D amorphous solids

Abstract
The underlying structural disorder renders the concept of topological defects in amorphous solids difficult to apply and hinders a first-principle identification of the microscopic carriers of plasticity and of the regions more prone to structural rearrangements (“soft spots”). Recently, it has been proposed that well-defined topological defects can still be identified in glasses, and correlated to local and global plasticity, by looking at the eigenvector field or the particle displacement field. Nevertheless, all the existing proposals and analyses are only valid in two spatial dimensions. In this work, we propose the idea of using hedgehog topological defects to characterize the plasticity of 3D glasses and to geometrically predict the location of their soft spots. We corroborate our proposal by simulating a Kremer-Grest 3D polymer glass, and by using both the normal mode eigenvector field and the displacement field around large plastic events. Our results confirm that a topological characterization of plasticity in glasses is feasible, and provide a concrete realization of this program in 3D amorphous systems.
Introduction – Topological defects, such as dislocations and disclinations, play a pivotal role for the description of mechanical failure in crystalline solids since they provide the microscopic origin of their plasticity Taylor (1934); Polanyi (1934); Orowan (1934). Topological defects also provide the building blocks for the theory of crystal melting in two dimensions Kosterlitz and Thouless (1973); Halperin and Nelson (1978); Young (1979). Despite several attempts Spaepen (1978); Cotterill (1979); Blackett (1979); Morris (1979); Popescu (1984); Egami and Vitek (1984); Shi (1993); Chaudhari et al. (1979); Egami et al. (1980); Acharya and Widom (2017); Nelson (1983); Steinhardt and Chaudhari (1981), topological defects associated to translational symmetry (dislocations and disclinations) cannot be formally defined within the structure of amorphous solids such as glasses, because of the absence of translational long-range order. This difficulty remains the main cause for the lack of a microscopic theory of plasticity and mechanical deformations in glasses Hufnagel et al. (2016); Rodney et al. (2011); Zaccone (2023).
More in general, the definition and identification of “defects” constitute the main obstacle for establishing a robust causal relation between structure and dynamics in amorphous solids (see e.g. Wang (2019); Cheng and Ma (2011)). So far, most of the approaches to solve this problem are still based on ad-hoc coarse-grained descriptions Nicolas et al. (2018), vibrational soft modes Rottler et al. (2014); Manning and Liu (2011); Ding et al. (2014); Chen et al. (2011); Kapteijns et al. (2020); Tanguy et al. (2010); Smessaert and Rottler (2014a), shear transformation zones revisited under the lens of anharmonic quasi-localized modes Zylberg et al. (2017); Richard et al. (2021); Gartner and Lerner (2016); Ruan et al. (2022); Xu et al. (2018, 2021), phenomenological structural indicators Richard et al. (2020); Widmer-Cooper et al. (2008) or machine-learning algorithms Fan and Ma (2021); Ciarella et al. (2023).
Two recent works Baggioli et al. (2021); Wu et al. (2023) (see also Baggioli (2023)) have shown that not only topological concepts can be applied (albeit with proper generalizations) to amorphous solids, as done for vortices in superfluids or dislocations in crystals, but that they can be also extremely useful to predict their mechanical properties both on a local and a global scale.
A first approach Baggioli et al. (2022) is based on the idea of continuous Burger vector Kleman and Friedel (2008) applied to the displacement field, and has been shown to be able to accurately locate plastic events under deformation and the yielding point Baggioli et al. (2021). It is likely that this concept is related to the geometric charges previously discussed as the carriers of plastic screening in glasses Lemaître et al. (2021). The second approach Wu et al. (2023) relies on the identification of vortex-like defects in the eigenvector field that strongly correlate with the plastic spots and display tight connections with the widely used concept of shear transformation zones Desmarchelier et al. (2024). This second type of vortex-like topological defects have been recently experimentally identified in a 2D colloidal glass Vaibhav et al. (2024), confirming the simulation results.
Importantly, all the proposals and analyses so far apply only to two-dimensional (2D) amorphous solids and lack direct application to more realistic 3D scenarios. In Cao et al. (2018), disclination-like defects, defined using polytetrahedral order Nelson and Spaepen (1989), were considered in a 3D experimental granular matter system and were shown to correlate with plasticity. In Bera et al. (2024), a 2D slicing method has been used to study vortex-like topological defects in a 3D simulation polymer glass revealing a strong clustering tendency of negative charged topological defects right before large plastic events.
As of today, a full-fledged definition of topological defects in 3D glasses and a direct proof of their connection to plasticity in three dimensions are still lacking. In this work, we present a solution to these problems.
Hedgehog topological defects – Given the large variety of topological defects that can be defined within a three-dimensional vector field (point defects, line defects and surface defects) Kleinert (1989); Nelson (2002), we suggest that hedgehog topological defects (HTD) are the simplest choice to start with. We adapt the definition of HTDs from nematic liquid crystals Kleman and Lavrentovich (2003, 2006); Hölbl et al. (2023) and Heisenberg magnets Fumeron and Berche (2023); Berg and Lüscher (1989); Holm and Janke (1994); Berg and Lüscher (1981) to the eigenvector and displacement fields of a 3D simulation glass.
For a generic 3D continuous vector field $\vec{s}$, we define the hedgehog topological charge as Kleman and Lavrentovich (2003, 2006); Hölbl et al. (2023)
[…]
where $\Sigma$ is an arbitrary closed 2D surface parametrized by coordinates $(x_{1},x_{2})$. $Q$ represents a topological invariant in the sense that it does not change under continuous and smooth deformations of the manifold.
Figure 1: (a) Discrete vector at a simple cubic lattice corners illustrating a radially outward field for a monopole with $Q=+1$ (red). A vertex $x^{*}$ is shown where $\vec{s}_{1}$, $\vec{s}_{2}$, $\vec{s}_{3}$ and $\vec{s}_{4}$ field at four corners of the vertex. (b) Continuous vector field (left) and discrete vector field (right) for a monopole. (c) Radially inward field representing an anti-monopole with $Q=-1$ (blue). (d) Continuous (left) and discrete (right) vector fields for $Q=-1$.
From our simulations, we first obtain a discrete vector field ${\vec{s}}_{l,m,n}$ at each lattice grid point $(l,m,n)$ in a simple cubic lattice. More details are provided in the Supplementary Material (SM). Let $\vec{s}_{1}$, $\vec{s}_{2}$, $\vec{s}_{3}$, and $\vec{s}_{4}$ be the unit vectors of the field associated with one face of the cube, where $\vec{s}_{1}$, $\vec{s}_{2}$, and $\vec{s}_{3}$ are positioned at the corners of triangle $i$ on that face, chosen such that the sequence $1-2-3-1$ corresponds to a counterclockwise rotation along the outward normal to the surface of the triangle. See Fig. 1(a) for a representation of these vectors on the example of a radially outgoing field (monopole). The topological charge $Q$ enclosed by the unit cube is then expressed in a discrete fashion as Holm and Janke (1994); Berg and Lüscher (1981),
$Q=\sum_{x^{*}}q(x^{*}),$ (2)
where the sum is carried over all $6$ vertices of the unit cube and $x^{*}$ is the vertex associated with the four vectors $\vec{s}_{1}$, $\vec{s}_{2}$, $\vec{s}_{3}$ and $\vec{s}_{4}$ (as shown in Fig. 1(a)). Here, $q(x^{*})$ is expressed as
$q(x^{*})=\frac{1}{4\pi}[(\sigma_{s}A)(\vec{s}_{1},\vec{s}_{2},\vec{s}_{3})+(\sigma_{s}A)(\vec{s}_{1},\vec{s}_{3},\vec{s}_{4})],$ (3)
where $(\sigma_{s}A)(\vec{s}_{1},\vec{s}_{2},\vec{s}_{3})$ is the signed area of the spherical (non-Euclidean) triangles with corners $\vec{s}_{1}$, $\vec{s}_{2}$ and $\vec{s}_{3}$. The sign of the area can be calculated as ${\sigma_{s}=\rm{sign}}\{\vec{s}_{1}\cdot[\vec{s}_{2}\times\vec{s}_{3}]\}$. Finally, the value of $A$ is calculated as Holm and Janke (1994); Berg and Lüscher (1981),
$\cos\left(\frac{A}{2}\right)=\frac{1+\vec{s}_{1}\cdot\vec{s}_{2}+\vec{s}_{2}\cdot\vec{s}_{3}+\vec{s}_{3}\cdot\vec{s}_{1}}{\sqrt{2(1+\vec{s}_{1}\cdot\vec{s}_{2})(1+\vec{s}_{2}\cdot\vec{s}_{3})(1+\vec{s}_{3}\cdot\vec{s}_{1})}}.$ (4)
A visual representation of this algorithm for the case of hedgehog defects with charges $Q=\pm 1$ (monopole and anti-monopole) is presented in Fig. 1.
3D polymer glass model – We simulate a 3D polymer glass system using the Kremer-Grest model Kremer and Grest (1990) with $N=10{,}000$ monomers of alternating masses of $m_{1}=1$ and $m_{2}=3$ Kriuchevskyi et al. (2020) interacting via the Lennard-Jones (LJ) potential ($u_{\text{LJ}}$) and the finite extensible nonlinear elastic (FENE) potential ($u_{\text{FENE}}$). We set the dihedral bending rigidity of the bonds to zero, as appropriate for flexible chains Palyulin et al. (2018). More details can be found in the SM.
We define the $3N\times 3N$ Hessian matrix, with elements written in terms of the total potential energy $u(r)=u_{\text{LJ}}(r)+u_{\text{FENE}}(r)$,
$H_{ij}^{\alpha\beta}=\frac{1}{\sqrt{m_{i}m_{j}}}\frac{\partial^{2}u(\vec{r})}{\partial{r}_{i}^{\alpha}\partial{r}_{j}^{\beta}}.$ (5)
Here, $i$ and $j$ are particle indices, while $\alpha$ and $\beta$ denote the spatial directions ($x$, $y$, $z$). Here $m_{i}$ represents the mass of the $i$-th particle, and $r_{i}^{\alpha}$ is the $\alpha$-component of the position vector of the $i$-th particle.
The eigenvalues $\lambda_{k}$ (or equivalently eigenfrequencies $\omega_{k}=\sqrt{\lambda_{k}}$) and their associated eigenvectors are obtained by direct diagonalization of the Hessian matrix. In the rest of the manuscript, we will use the symbols $\omega_{k}$ and $\omega$ interchangeably.
Soft spots from the eigenvector field – We start by applying our algorithm to the 3D eigenvector field of the simulated polymer glass system. To determine the HTDs, we derived the smoothed eigenvector field on a simple cubic lattice (see SM for details). For direct visualization, we present a snapshot of the smooth eigenvector field that exhibits an hedgehog anti-monopole and a monopole in the SM, Fig. S2.
By scanning the eigenvector field at different frequencies, we have verified that the total number of HTDs scales at low frequency (up to $\omega\approx 4$) as $N_{\text{HTD}}\propto\omega^{2}$ (shown in SM, Fig. S3(b)) and correlates with the corresponding density of states, which exhibits Debye scaling, $D(\omega)\propto\omega^{2}$ (see SM, Fig. S3(a)) at low frequency. This result is consistent with previous findings in 2D simulation and experimental works Wu et al. (2023); Vaibhav et al. (2024).
After confirming the validity of the 3D algorithm presented above, we investigate the physical properties of the HTDs, and in particular whether they bear any connection with plastic soft spots. In order to do so, we follow Smessaert and Rottler (2014b) and we identify the soft spots in the 3D model glass system using the softness parameter
$\phi_{i}=\frac{1}{N_{m}}\sum_{j=1}^{N_{m}}\frac{|{\bf e}^{(i)}_{j}|^{2}}{m_{i}\omega_{j}^{2}},$ (6)
where $N_{m}$ is the number of low frequency modes used in the analysis and ${\bf e}^{(i)}_{j}$ is the eigenvector of the $j$-th mode, corresponding to the $i$-th particle. To define $N_{m}$, we have considered modes up to $\omega=5$. Finally, the plastic soft spots are identified using the locations of the $5\%$ of particles having higher values of $\phi_{i}$.
Figure 2: (a) The color map visualizes the 3D average topological charge density. The black symbols are the locations of soft spots, identified using the softness parameter, Eq. (6). (b)-(f) Pair correlation functions between HTDs: monopoles (p) and anti-monopoles (n), and plastic soft spots (s) for different frequencies. Note that the curves for $\omega>1.1$ are shifted upward by constant factors for better visualization in this and all other similar data sets. See SM, Fig. S4, for extended data.
In Fig. 2(a), we show a 3D color map indicating the intensity of the average topological charge density (see SM for details). Red and blue colored regions represent therefore areas with higher concentrations of positive and negative topological charge $Q$, respectively. In the same box, we provide the locations of the plastic soft spots using black symbols. In order to test the correlation between HTDs and plasticity, in Fig. 2(b)-(g), we present various pair correlation functions between HTDs (p), anti-monopole HTDs (n) and plastic soft spots (s) for different frequencies.
In Fig. 2(b)-(c), the correlation between HTDs and soft spots is shown. Our results reveal that both positive and negative defects exhibit strong correlations with soft spots at short and intermediate distances at low frequency, that nevertheless vanish quickly by increasing $\omega$ (see Fig. S4 in the SM). This outcome presents marked differences with the 2D case Wu et al. (2023), where soft spots predominantly correlate with negative defects at short distances. This might be attributed to the different nature of point defects in 2D and 3D systems Kleman and Lavrentovich (2003).
Moreover, HTDs with same charge also display a peak in their spatial correlation at approximately the same distance, see Fig. 2(d) and Fig. 2(f), indicating a widespread tendency to clustering of HTDs. This may be explained with the fact that HTDs tend to nucleate more in regions of significant nonaffine motions Baggioli et al. (2022), which are spatially localized DiDonna and Lubensky (2005). Hence, it is statistically more likely to have charges, even with same sign, next to each other. Finally, in Fig. 2(e), we observe a very strong short-distance correlation between monopole and anti-monopole defects that suggests that defects with opposite charges attract each other and tend to pair together to minimize the local topological charge. This correlation is very strong at low frequency and becomes weaker at higher $\omega$.
Predicting plasticity from the displacement vector – Inspired by the proposal of looking at topological defects in the dynamical displacement field $\vec{u}$ Baggioli et al. (2021), rather than in the eigenvector field, we now apply the same 3D algorithm to find HTDs in $\vec{u}$. We notice that there is no direct relation between the displacement field and a single eigenvector $\mathbf{e}_{k}$, despite $\vec{u}$ can be decomposed onto the complete basis set of eigenvectors. Therefore, a correlation between topological defects in the eigenvector field and in the displacement field is not obvious. Nevertheless, around plastic events, it is known that the dynamical response is dominated by a few localized and low frequency modes (see e.g. Yang et al. (2020)), providing a possible connection between the topology of the low-frequency eigenvectors and that of the displacement field.
In our study, we apply quasi-static shear to our system using the athermal quasi static (AQS) protocol Lemaître and Maloney (2006) with strain step $\delta\gamma=10^{-3}$. The stress-strain curve is presented in the SM (Fig. S1(c)) and displays a yielding instability at $\gamma^{*}\approx 0.099$. For presentation purposes, here we focus on four plastic events, including the yielding point, and consider the non-affine displacement field at these specific strain values. In order to test the universality of our findings, we define the soft spots using two different methods: (I) the softness parameter defined in Eq. (6) (and used in the previous analysis) and (II) the value of the particle level non-affine displacement field $\vec{u}_{\text{NA}}$. Within this second method, we identify the plastic soft spots by looking at the locations of the $5\%$ of particles having higher values of $|\vec{u}_{\text{NA}}|$.
Figure 3: (a)-(d) Spatial correlation functions between soft spots (identified using the softness parameter, Eq. (6)) and HTDs within the non-affine displacement field across four plastic events at different $\gamma$. Correlations between soft spots and HTDs with $Q=+1$($-1$) are shown in red (blue).
In order to test the predictability of the plastic regions from pure static configurations, we obtain the soft spots using method (I) applied just before the plastic events. In Fig. 3(a)-(d) we show the spatial correlation between HTDs within non-affine displacement field (see SM for details) and the soft spots identified with the softness parameter $\phi_{i}$ (method I) for four plastic events at $\gamma=0.099,~{}0.120,~{}0.126$ and $0.154$ (see Fig. S1(c) for associated stress drop). From there, it is evident that both positive and negative HTDs with displacement field display a clear correlation with the soft spots that is rather localized around $r\approx 1$. Once again, no marked distinction between negative and positive defects is observed.
[…]
In order to confirm this result, in the bottom panels of Fig.4, we show the spatial correlation function between monopoles in the displacement field and soft spots (red data) and between anti-monopoles in the displacement field and soft spots (blue data). For all the three values of strain considered, we observe a very strong spatial correlation between the HTDs in the displacement field and the soft spots, with a pronounced peak appearing at $r\approx 1$. Interestingly, the distance at which this correlation emerges is compatible with the low-frequency results presented in Fig. 2, where the HTDs have been identified in the eigenvector field, and the results in Fig. 3 as well. This confirms our initial speculation that, around plastic events, the topology of the low-frequency eigenvectors, and even its spatial characteristics, are compatible, and in fact very similar, to those observed in the displacement vector.
Finally, we notice that there is no apparent difference between negative ($Q=-1$) and positive ($Q=+1$) HTDs, or at least in their correlation with plasticity. This is independent of the methods employed to define the HTDs and also the soft spots, and therefore rather universal. Interestingly, this represents a dissimilarity with respect to the analysis using the eigenvector field in 2D (or in the 2D-slicing method for 3D systems), where negative defects seem to have a privileged role Wu et al. (2023); Desmarchelier et al. (2024); Bera et al. (2024). In addition, we notice that the results shown in Fig. 4, using the non-affine displacement field, display a rather long-range structure in which spatial correlations survive up to large distances, $r\approx 5$. This very long correlation length might depend on several parameters and deserves further investigation. Moreover, the spatial structure of the HTDs and the spatial decay of their correlation with soft spots might provide useful insights to achieve a microscopic definition of the elastic screening length proposed in other macroscopic approaches to amorphous solids Lemaître et al. (2021).
Figure 4: (a)-(c) The location of HTDs with charges $Q=+1$ and $Q=-1$ in the displacement vector field at different $\gamma$ are shown in red and colorblueblue, respectively. The yellow regions are the plastic soft spots, identified using the softness parameter, Eq. (6). (d)-(f) The spatial correlation between between monopoles (in the displacement field) and soft spots (red data) and anti-monopoles (in the displacement field) and soft spots (blue data) for the same values of the strain $\gamma$. For this analysis, soft spots were identified from the top $10\%$ of highest $|\vec{u}_{\text{NA}}|$ values.
Conclusions – In this work, we proposed to use point-like hedgehog topological defects (HDTs) to characterize the plasticity of 3D amorphous solids and to locate their soft spots using well-defined mathematical concepts. First, borrowing from the physics of liquid crystals Kleman and Lavrentovich (2003, 2006); Hölbl et al. (2023) and Heisenberg magnets Fumeron and Berche (2023); Berg and Lüscher (1989); Holm and Janke (1994); Berg and Lüscher (1981), we have outlined a concrete and practical procedure to identify HDTs in 3D glasses – a method that is suitable for experimental discrete data sets as well.
We have then applied this proposed algorithm to both the eigenvector field and the displacement vector of a simulated 3D polymer glass model and verified the feasibility of this method. For the HTDs identified in the eigenvector field, we have proved that the number of HTDs as a function of frequency follows the same trend as in previous 2D studies, and directly correlates with the vibrational density of states at low frequency. We have also shown a strong tendency in forming oppositely charged monopole/anti-monopole pairs of topological defects and even clusters with larger number of defects. Moreover, we have revealed a very strong spatial correlation between the hedgehog topological defects and the plastic soft spots at short distances.
Finally, we considered HTDs in the 3D displacement vector field for a quasi-static shear deformation and for different large plastic events, including the yielding point. In this analysis, we have found that HTDs correlate very strongly with the extended plastic soft spots from which the plastic instability originates, that are defined using two different methods based respectively on the softness parameter Smessaert and Rottler (2014b) and the amplitude of the non-affine displacement field. This result proves that the location of HTDS, and their yet undisclosed dynamics, might play a pivotal role to achieve a microscopic understanding of plasticity and yielding in 3D glasses.
Interestingly, the analysis of topological defects in 3D amorphous solids revealed that the correlation with the plastic soft spots is independent of the charge of the defect, unlike previous findings in 2D systems Wu et al. (2023); Desmarchelier et al. (2024); Bera et al. (2024) and in the 2D-slicing of 3D systems Bera et al. (2024). At present, we do not have an explanation for this intriguing phenomenon that requires further study.
Before concluding, we also notice that the correlation between plasticity and topological defects appears stronger when the latter are defined using the displacement field, rather than the low-frequency eigenvector fields. This is promising since, apart from few exceptions (e.g., Vaibhav et al. (2024)), the eigenvector field is not a quantity directly accessible via experiments, whereas the displacement vector is.
On a general ground, our work proves that the idea of using geometric topology to describe, predict and rationalize plastic properties of amorphous solids can be successfully extended to three dimensional systems, such as real materials. Aside from the theoretical part of this development, this result provides a formidable occasion to test these ideas on realistic experimental amorphous solids for which the method we proposed can be directly applied. For example, three dimensional sheared granular systems Cao et al. (2018) constitute a possible candidate platform to test these ideas.
Acknowledgments – M.B. thanks W. Kob, Z. Wu, J. Zhang, Y. Wang, C. Jiang, Z. Zheng, H. Tong for useful discussions. We would like to thank A. Liu and T. Petersen for ongoing collaborations and endless discussions about this topic as well. M.B. acknowledges the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01) and the sponsorship from the Yangyang Development Fund. A.Z. gratefully acknowledges funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”. A.Z. gratefully acknowledges the Niedersächsische Akademie der Wissenschaften zu Göttingen in the frame of the Gauss Professorship program. A.Z. and A.B. gratefully acknowledge funding from US Army Research Office through Contract No. W911NF-22-2-0256.
Supplementary Material
In this Supplementary Material (SM), we provide further details about the numerical and analytical computations presented in the main text. Moreover, we show further analysis to corroborate our findings.
Model and numerical methods
In this study, we implement the Kremer-Grest model Kremer and Grest (1990) to simulate a coarse-grained polymer system. This system comprises linear polymer chains, each consisting of 50 monomers, with alternating monomer masses of $m_{1}=1$ and $m_{2}=3$ Kriuchevskyi et al. (2020). The total number of monomers in our system was $N=10{,}000$.
Figure S1: (a) Schematic representation of the polymer chains with alternating masses $m_{1}$ and $m_{2}$. The FENE bonds and the LJ interactions range $r_{c}$ from a monomer (A) are shown. (b) A typical simulation configuration is presented where monomer breads with masses $m_{1}$ and $m_{2}$ are indicated in red and blue, respectively and FENE bonds by green colors. (c) Stress versus strain curve. The yielding point is indicated by the dashed black line at $\gamma=0.099$.
The interactions between monomers in our model are governed by the truncated and shifted Lennard-Jones (LJ) potential:
$u_{\text{LJ}}(r)=\begin{cases}V(r)-V(r_{c}),&\text{for}~{}r<r_{c}=2.5\sigma,\\ 0,&\text{otherwise.}\end{cases}$ (S1)
Here, $V(r)$ represents the standard Lennard-Jones potential:
$V(r)=4{\varepsilon}\left[\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right],$ (S2)
where $\varepsilon$ is the interaction strength, and $\sigma$ is the monomer diameter. Here $r_{c}=2.5\sigma$ is the interaction distance.
In addition to LJ interactions, covalent bonds along the polymer chains are modeled using the finite extensible nonlinear elastic (FENE) potential:
$u_{\text{FENE}}(r)=-\frac{Kr_{0}^{2}}{2}\ln\left[1-\left(\frac{r}{r_{0}}\right)^{2}\right].$ (S3)
We set $K=30$ and $r_{0}=1.5\sigma$ in our simulations. In Fig. S1(a) we schematically show these interactions. The polymer chains were simulated within a cubic box with periodic boundary conditions applied in all three dimensions. The system was equilibrated using the LAMMPS simulation package Plimpton (1995). We show a typical configuration in Fig. S1(b). Following equilibration, we employed an athermal quasi-static (AQS) deformation protocol. This involved quenching the glass sample to absolute zero temperature and subsequently applying quasi-static shear with a strain increment of $\delta\gamma_{xz}=0.001$. The units of mass, length, and time of our simulation are taken as $m_{1}$, $\sigma$, and $\tau=\sqrt{m_{1}\sigma^{2}/{\varepsilon}}$, respectively. We set $m_{1}$, $\sigma$, and $\varepsilon$ to unity. Throughout our simulations, polymer chains were considered fully flexible, with no angle-bending terms included in the potential. The stress-strain curve in the shear plane is presented in Fig. S1(c).
Calculation of smoothed eigenvector field
Given the Hessian matrix, we compute the eigenvector field ${\mathbf{e}}_{k}=(e^{x}_{k},e^{y}_{k},e^{z}_{k})$ for each eigenmode with eigenfrequency $\omega_{k}$ ($k=1,2,\ldots,3N$). This eigenvector field is interpolated onto a $40\times 40\times 40$ simple cubic lattice with grid spacing of $\sigma/2$ in all directions, superimposed on the 3D simulation box. The eigenvector field at each lattice site $\vec{r}=(l,m,n)$ is then obtained as Wu et al. (2023)
$e^{\mu}_{k}(\vec{r})=\frac{\sum_{d_{i}<R_{\text{cut}}}w(\vec{r}-\vec{r}_{i})e^{\mu}_{i}}{\sum_{d_{i}<R_{\text{cut}}}w(\vec{r}-\vec{r}_{i})},\quad\mu=x,y,z,$ (S4)
where $\vec{r}_{i}$ is the position of particle $i$, and $w(\vec{r}-\vec{r}_{i})$ is a Gaussian weight function given by $w(\vec{r}-\vec{r}_{i})=\exp(-|\vec{r}-\vec{r}_{i}|^{2}/r_{c}^{2})$ with $r_{c}=1$. The distance $d_{i}=|\vec{r}-\vec{r}_{i}|$ is between the lattice site and the particle $i$, and $R_{\text{cut}}=4\sigma$ is the cut-off distance for interpolation.
Figure S2: Normalized eigenvector field for a portion of the simulation box exhibiting a monopole defect (red) and an anti-monopole one (blue). These data correspond to $\omega=3$.
Calculation of topological charge density
In order to compute the topological charge density, we divide the entire 3D simulation box into cubical cells of side $\sigma/2$. Within each such cell we count number of HTDs $N^{+}$ and $N^{-}$ with $Q=+1$ and $Q=-1$, respectively, for $\omega_{k}<4$. The topological charge density inside each cubical cell is defined as
$\bar{C}=\frac{\sum_{\omega_{k}}(N^{+}-N^{-})/{\omega}_{k}^{2}}{\sum_{\omega_{k}}1/{\omega}_{k}^{2}},$ (S5)
where the weight factor is motivated by the quadratic dependence of the number of HTDs at low frequency.
Number of HTDs and vibrational density of states
In Fig. S2, we show a snapshot of a 3D eigenvector field presented two clear topological defects with opposite charge.
In Fig. S3(a), we present the vibrational density of states (vDOS) of the polymer glass model at $T=0$. As highlighted in the inset, at low frequency the vDOS exhibits a distinct Debye scaling, $D(\omega)\propto\omega^{2}$. In panel (b) of the same figure, we present the number of hedgehog topological defects in the eigenvector field $N_{\text{HTD}}$ as a function of frequency $\omega$. At low frequency, the number of defects also scales as $\omega^{2}$, implying a correlation with the vDOS.
Figure S3: (a) The vibrational density of states $D(\omega)$ of the polymer glass model at $T=0$ extracted from the direct diagonalization of the Hessian matrix. The inset highlights the Debye scaling of the vDOS at low frequency. (b) The number of hedgehog topological defects in the eigenvector field $N_{\text{HTD}}$ as a function of frequency $\omega$. The inset zooms on the low frequency region and highlights the $\omega^{2}$ scaling.
Calculation and interpolation of non-affine displacement field on a cubic lattice
We derive the displacement field as Baggioli et al. (2021); Bera et al. (2024) $\vec{u}_{i}={\vec{r_{i}}}(\gamma)-{\vec{r}_{i}}(\gamma-\delta\gamma)$, where $\vec{r}_{i}(\gamma)$ and $\vec{r}_{i}(\gamma-\delta\gamma)$ represents the locations of the particle $i$ at strains $\gamma$ and $\gamma-\delta\gamma$, respectively. The affine component for the simple shear scenario is calculated as $\vec{u}_{A}=z\delta\gamma\hat{x}$. Finally we obtain the non-affine displacement field by subtracting the affine component from the total displacement, $\vec{u}_{\text{NA}}=\vec{u}_{i}-\vec{u}_{A}$ Baggioli et al. (2021); Bera et al. (2024). To obtain HTDs within the non-affine displacement field, we interpolated the discrete non-affine displacement field $\vec{u}_{\text{NA}}$ onto a simple cubic lattice grid with a side length $\sigma/2$ using the standard Python package griddata with linear interpolation.
Appendix A Extended data and correlation between soft spots identified with different methods
In Fig. S4, we provide extended data for the spatial correlation between between soft spots (s) identified using the softness parameter, Eq. (6), and negative (n) and positive (p) HTDs. Low-frequency data are provided in Fig. 2 in the main text. From Fig. S4, the vanishing of the correlation at high frequency ($\omega\gtrapprox 2$) is evident.
Finally, for completeness, in Fig. S5 we show the spatial correlation between the plastic soft spots identified with method I (softness parameter, Eq. (6)) and method II (non-affine displacement field). From there, we observe that the two definitions provide a strong correlation at short distance that, nevertheless, does not survive for large length scales.
Figure S4: Extended data showing the spatial correlation between soft spots (s) identified using the softness parameter, Eq. (6), and negative (n) and positive (p) HTDs for several higher frequencies. See Fig. 2 in the main text for low-frequency data.
Figure S5: Spatial correlation between soft spots identified using the softness parameter, Eq. (6), and the non-affine displacement field. Different lines correspond to different values of strain, located around major plastic events along the stress-strain curve, Fig. S1(c).